# Foreground Removal in Ground-Based CMB Observations Using a Transformer Model

### v1

@@ 5 Conclusions
We thank Yi-Ming Wang, Yong-Ping Li, Hua Zhai, Wen-Zheng Chen for useful discussion. This study is supported by the Postdoctoral Fellowship Program of CPSF under Grant Number GZC20241729, the China Postdoctoral Science Foundation under Grant Number 2024M763267, the National Nature Science Foundation of China under grant Nos. 12403005 12473004, and 12021003, the National Key R&D Program of China No. 2020YFC2201601 and 2020YFC2201603, and the Fundamental Research Funds for the Central Universities. Some of the results in this paper have been derived using the HEALPix (page: https://healpix.sourceforge.io/). We would like to express our gratitude for the following software packages: Healpy, CAMB, NaMaster, LensPyx, PySM, and PySM3.

### v2

@@ 5 Conclusions
We thank Yi-Ming Wang, Yong-Ping Li, Hua Zhai, Wen-Zheng Chen for useful discussion. This study is supported by the Postdoctoral Fellowship Program of CPSF under Grant Number GZC20241729, the China Postdoctoral Science Foundation under Grant Number 2024M763267, the National Nature Science Foundation of China under grant Nos. 12403005 12473004, and 12021003, the National Key R&D Program of China No. 2020YFC2201601 and 2020YFC2201603, and nhe Fundamental Research Funds for the Central Universities. Some of the results in this paper have been derived using the HEALPix (page: https://healpix.sourceforge.io/). We would like to express our gratitude for the following software packages: Healpy, CAMB, NaMaster, LensPyx, PySM, and PySM3.